OpenStudy (anonymous):

MEDALS A triangle has two sides of lengths 4 and 15. What value could the length of the third side be? Check all that apply. A. 13 B. 11 C. 12 D. 4 E. 15 F. 19

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@briannaearls12 @Jumperman

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

B

OpenStudy (jumperman):

What kind of triangle is it?

OpenStudy (anonymous):

15-4=11

OpenStudy (igreen):

The Pythagorean Theorem \(a^2 + b^2 = c^2\) Where \(a\) and \(b\) are the two side lengths and \(c\) is the hypotenuse.

OpenStudy (anonymous):

Does it give a specific triangle? Like acute, obtuse, or equilateral?

OpenStudy (anonymous):

no

OpenStudy (bass53):

B

OpenStudy (anonymous):

15 - 4 = 11 - The sides have to connect so your answer is B.

OpenStudy (anonymous):

it has nothing to do with pythagoras

OpenStudy (anonymous):

you have to check that the sum of any two sides is bigger than the third side